I'm stumped and i can't seem to get this. Directions say
Find the first five terms of the given sequence
u0=1
u1=1 and and un =(n)Un-1 for n(is greater than or equal to) 2

Answer by stanbon(75887)

Find the first five terms of the given sequence
u0=1
u1=1 and and un =(n)U(n-1) for n(is greater than or equal to) 2
-------------
u(0) = 1
u(1) = 1
u(2) = 2*1 = 2
u(3) = 3*2 = 6
u(4) = 4*6 = 24
u(5) = 5*24 = 120
u(6) = 6*120 = 720
========================
Cheers,
Stan H.